• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name brads

    Log entry time 17:10:31 on May17,2008

    Entry number 238455

    keyword=BigBite magent current changes since 16:30

    Shift workers: It is best if you halog changes to the experimental setup at the time they are made. If you're doing a scan, then log a summary/conclusions at the end.

    Also, the BB current in the CODA runstart log entries is not being kept up to date with these changes...

    Changed to 400A at 16:38.

    Changed to 200A at 16:42.

    Changed to 100A at 17:05.


    FIGURE 1

    FIGURE 2